# ORACLE: Explaining Feature Interactions in Neural Networks with ANOVA

### v3

@@ Abstract
We introduce ORACLE, a framework for explaining neural networks on tabular data and scientific factorial designs. ORACLE summarizes a trained network’s prediction surface with main effects and pairwise interactions by treating the network as a black-box response, discretizing the inputs onto a grid, and fitting an orthogonal factorial (ANOVA-style) surrogate—the $L^{2}$ orthogonal projection of the model response onto a finite-dimensional factorial subspace. A simple centering and $\mu$-rebalancing step then expresses this surrogate as main- and interaction-effect tables that remain faithful to the original model in the $L^{2}$ sense. The resulting grid-based interaction maps are easy to visualize, comparable across backbones, and directly aligned with classical design-of-experiments practice. On synthetic factorial benchmarks and low- to medium-dimensional tabular regression tasks, ORACLE more accurately recovers ground-truth interaction structure and hotspots than Monte Carlo SHAP-family interaction methods, as measured by ranking, localization, and cross-backbone stability. In latent image and text settings, ORACLE clarifies its scope: grid-based factorial surrogates are most effective when features admit an interpretable factorial structure, making ORACLE particularly well-suited to scientific and engineering workflows that require stable, DoE-style interaction summaries.
@@ 2 Related Work
Recent work on global interaction explanations has largely built on the Shapley framework. Faith-Shap introduces an axiomatic Shapley interaction index as a canonical notion of global interaction strength (tsai2023faith), while SHAP-IQ, SVARM-IQ, and KernelSHAP-IQ provide Monte Carlo and kernel-based schemes to efficiently approximate any-order Shapley interactions in practice (fumagalli2023shap; kolpaczki2024svarm; fumagalli2024kernelshap). These algorithms and variants are consolidated and benchmarked in the shapiq library, which compares Shapley-based interaction estimators across models and datasets (muschalik2024shapiq). Complementary work analyzes how global feature effects decompose into main and interaction components (herbinger2024decomposing) and uses Shapley interactions to probe structured phenomena such as syntax and speech representations in large models (singhvi2024knowing). However, these approaches mostly return scalar interaction indices rather than explicit pairwise interaction maps, and are rarely evaluated against an interaction oracle or for stability across backbones.
@@
Functional ANOVA has re-emerged as a central tool for interpretable machine learning, motivating intrinsically interpretable models that decompose predictions into main effects and low-order interactions, as well as neural architectures that directly parameterize functional-ANOVA components or screen interactions before fitting a factorial surrogate, enabling higher-order terms with statistical guarantees (hu2025interpretable; kohler2025achieving; choi2025meta; park2025tensor). In engineering design and global sensitivity analysis, interpretable neural networks are combined with functional-ANOVA decompositions and grouped sensitivity indices to study complex physical interactions from observational data (dolar2024data). Complementary work in design-of-experiments couples factorial and fractional-factorial designs with modern machine learning to estimate main and interaction effects efficiently in industrial settings (fontana2023design; zubair2024coupling). Our work adopts this ANOVA/DoE perspective but targets pairwise interaction maps of a fixed neural backbone and benchmarks interaction explanation methods, rather than proposing another globally additive predictive model.
@@
A complementary line of work develops 2D visualizations of how feature pairs jointly affect predictions. Regional effect plots and their extensions, including DALE and recent fast pipelines, efficiently generate ALE/PDP-style 2D regional plots across many features and models (herbinger2022repid; gkolemis2023dale; gkolemis2024fast). The vivid framework and PDPilot further support heatmaps and interactive analytics for ranking and exploring 1D/2D effect patterns (inglis2022visualizing; kerrigan2025pdpilot), while toolkits such as Effector integrate these regional explanations into Python workflows (gkolemis2024effector). Beyond these plotting systems, transformed-feature-space approaches summarize joint and interaction effects under strong predictor dependence, yielding alternative 2D visualizations that better respect data geometry (brenning2023interpreting).
@@
A growing line of work studies how sensitive explanations are to perturbations of inputs, parameters, or training conditions, and how consistent they remain across independently trained models. (agarwal2022rethinking) formalize stability metrics for attribution methods and show that popular saliency techniques can be highly unstable, while (xue2023stability) obtain formal stability guarantees by smoothing the predictor. Algorithmic approaches include training objectives that directly regularize explanation stability (chen2024training) and VAE-based schemes that generate more stable LIME-style surrogates (xiang2023stable). Beyond single models, (watson2022agree) show that identical architectures with different random seeds can yield markedly different explanations, (wu2024improving) propose an explanation-consistency score between original and adversarial images, and (rai8evaluating) quantify consistency of LIME, SHAP, and PDP explanations under varying class-imbalance regimes. These works primarily target local attributions or saliency maps and study stability under perturbations or retraining, whereas we focus on the consistency of pairwise interaction rankings and interaction maps across independently trained backbones, measured via cross-backbone ranking and surface-alignment metrics.
@@
Several recent works explicitly evaluate how well interaction explanations recover known or controlled dependency structure. Examples include grammar-derived dependency relations with rank-based metrics in formal-language classifiers (jumelet2023feature), Rashomon-set analyses with interaction scores and Halo plots (li2023exploring), error-controlled discovery procedures with FDR guarantees on synthetic and semi-synthetic tasks (chen2025error), and stability-aware interaction evaluation on industrial intrusion-detection benchmarks (bouke2024application). XAI-Units and SPEX provide unit-test-style interaction benchmarks and LLM-scale setups with synthetic triggers and downstream QA tasks (lee2025xai; kang2025spex), while complementary work in NLP proposes unified evaluation frameworks for feature explanations with multiple faithfulness and robustness metrics (sun2025evaluating). In contrast, we develop a targeted evaluation protocol and metrics for global pairwise interaction maps in regression backbones.
@@ 3.2 Target: interaction map of a backbone
$g_{jk}(x_{j},x_{k})$ $=\mathbb{E}\bigl[f(X)\mid X_{j}=x_{j},X_{k}=x_{k}\bigr]-m_{j}(x_{j})-m_{k}(x_{k})-\mu$
@@ 3.3 Discretized factorial surrogate
We discretize each feature into $L$ bins, inducing a regular grid over $\mathcal{X}$. For feature $j$, let $\mathcal{B}_{j}=\{1,\dots,L\}$ index bins, and let $q_{j}(x_{j})\in\mathcal{B}_{j}$ denote the bin assignment for value $x_{j}$ (e.g., via empirical quantiles or uniform spacing over the observed range). A grid point is identified by a level vector $\ell=(\ell_{1},\dots,\ell_{d})$ with $\ell_{j}\in\mathcal{B}_{j}$, and we write $f(\ell)$ as shorthand for the backbone prediction at a representative input from that cell.
On this grid, we fit a linear factorial surrogate
$\hat{f}_{L}(\ell)\;=\;\hat{\mu}+\sum_{j=1}^{d}\hat{m}_{j}(\ell_{j})+\sum_{1\leq j<k\leq d}\hat{g}_{jk}(\ell_{j},\ell_{k}),$
where $\hat{m}_{j}:\mathcal{B}_{j}\to\mathbb{R}$ and $\hat{g}_{jk}:\mathcal{B}_{j}\times\mathcal{B}_{k}\to\mathbb{R}$ are discrete main-effect and interaction effect tables. In matrix form, this corresponds to a least-squares fit of a factorial design matrix $\Phi\in\mathbb{R}^{n\times p}$, whose columns encode cell indicators and contrast-coded factors, to the backbone responses $\{f(x_{i})\}$. Our framework, ORACLE, solves
$\hat{\beta}\;=\;\arg\min_{\beta}\frac{1}{n}\sum_{i=1}^{n}\bigl(f(x_{i})-\langle\beta,\phi(q(x_{i}))\rangle\bigr)^{2},$
where $\phi(q(x_{i}))$ denotes the basis evaluated at the discretized levels $q(x_{i})=(q_{1}(x_{i1}),\dots,q_{d}(x_{id}))$. The coefficients $\hat{\beta}$ can then be reshaped into the tables $\hat{\mu}$, $\{\hat{m}_{j}\}$, and $\{\hat{g}_{jk}\}$.
@@ 3.4 Recentering and $\mu$-rebalancing (identifiability)
A naive least-squares fit does not necessarily satisfy the standard identifiability constraints for factorial effect tables, such as zero-mean main effects and zero row/column sums for interaction tables. To obtain a well-defined decomposition, we apply a lightweight post-processing step that recenters the estimated tables and rebalances the overall mean.
For each feature $j$, we subtract the average of the main-effect table to enforce
$\sum_{\ell_{j}\in\mathcal{B}_{j}}\hat{m}_{j}(\ell_{j})\;=\;0.$
For each pair $(j,k)$, we adjust the interaction table $\hat{g}_{jk}$ so that its row and column sums vanish,
$\sum_{\ell_{j}\in\mathcal{B}_{j}}\hat{g}_{jk}(\ell_{j},\ell_{k})=0,\quad\sum_{\ell_{k}\in\mathcal{B}_{k}}\hat{g}_{jk}(\ell_{j},\ell_{k})=0\quad\text{for all }\ell_{k}\in\mathcal{B}_{k},\ \ell_{j}\in\mathcal{B}_{j},$
while preserving the overall mean on the grid. Finally, any constant offset induced by these centering operations is absorbed into $\hat{\mu}$ ($\mu$-rebalancing”). After this procedure, the discrete surrogate satisfies
$\hat{f}_{L}(\ell)\;=\;\hat{\mu}+\sum_{j=1}^{d}\hat{m}_{j}(\ell_{j})+\sum_{1\leq j<k\leq d}\hat{g}_{jk}(\ell_{j},\ell_{k}),$
with the usual sum-to-zero constraints at the discretized-grid level. This makes the effect tables interpretable as discrete approximations of the corresponding functional-ANOVA components of $f$ under $P_{X}$. We summarize the recentering and $\mu$-rebalancing procedure in Algorithm 1 (Appendix B).
@@ 3.5 Extracting interaction maps and strengths
The post-processed interaction tables $\hat{g}_{jk}:\mathcal{B}_{j}\times\mathcal{B}_{k}\to\mathbb{R}$ form discrete interaction maps for each feature pair $(j,k)$. These maps can be visualized directly as heatmaps over the $(X_{j},X_{k})$ grid, revealing regions where the backbone exhibits strong non-additivity. To summarize interaction strength in a single scalar, we compute the root-mean-square (RMS) over the grid,
$\hat{S}_{jk}\;=\;\Bigg(\frac{1}{|\mathcal{B}_{j}||\mathcal{B}_{k}|}\sum_{\ell_{j}\in\mathcal{B}_{j}}\sum_{\ell_{k}\in\mathcal{B}_{k}}\hat{g}_{jk}(\ell_{j},\ell_{k})^{2}\Bigg)^{1/2},$
and use $\{\hat{S}_{jk}\}_{j<k}$ to rank interactions. In our experiments, these estimated interaction maps and strengths serve as the global explanations produced by ORACLE. They are evaluated against oracle interaction maps derived from the backbone and $P_{X}$, and compared to SHAP-family baselines in terms of interaction selection, hotspot localization, cross-backbone transfer, scale alignment, and intervention-oriented utility.
@@ 4.1 Oracle interaction decomposition via functional ANOVA
$g_{jk}(x_{j},x_{k})$ $=\mathbb{E}\bigl[f(X)\mid X_{j}=x_{j},X_{k}=x_{k}\bigr]-m_{j}(x_{j})-m_{k}(x_{k})-\mu,$
and this representation coincides with the orthogonal decomposition in Definition 4.1 (see Appendix for details).
@@ 4.3 Consistency of interaction strengths and Top-$K$ selection
We now study the behavior of interaction strengths induced by the discrete interaction maps. For a fixed grid size $L$, define the discrete interaction strength
$S^{L}_{jk}=\Big(\mathbb{E}\big[g^{L}_{jk}(Z_{j},Z_{k})^{2}\big]\Big)^{1/2},$
and its empirical counterpart
$\hat{S}_{jk}=\Bigg(\frac{1}{|\mathcal{B}_{j}||\mathcal{B}_{k}|}\sum_{\ell_{j}\in\mathcal{B}_{j}}\sum_{\ell_{k}\in\mathcal{B}_{k}}\hat{g}_{jk}(\ell_{j},\ell_{k})^{2}\Bigg)^{1/2}.$
@@
Under the conditions of Proposition 4.6, we have
$\hat{S}_{jk}\xrightarrow{p}S^{L}_{jk}\quad\text{for all }1\leq j<k\leq d\quad\text{as }n\to\infty.$
If, in addition, the oracle maps $g_{jk}$ are square-integrable and sufficiently regular, then for each $(j,k)$,
$S^{L}_{jk}\to S^{\star}_{jk}\quad\text{as }L\to\infty.$
We next translate this convergence into guarantees for ranking and NDCG-style metrics. Let $S^{\star}_{(1)}\geq\dots\geq S^{\star}_{(M)}$ denote the sorted oracle strengths over all $M=\binom{d}{2}$ pairs, and similarly let $\hat{S}_{(1)}\geq\dots\geq\hat{S}_{(M)}$ be the sorted estimated strengths.
@@
for some fixed $K\in\{1,\dots,M-1\}$. Then, as $n\to\infty$ and $L\to\infty$,
$\mathbb{P}\big(\mathrm{Top}\text{-}K(\hat{S}_{jk})=\mathrm{Top}\text{-}K(S^{\star}_{jk})\big)\;\longrightarrow\;1,$
where $\mathrm{Top}\text{-}K(\cdot)$ denotes the set of indices of the $K$ largest entries.
@@
Under the assumptions of Theorem 4.8, the normalized discounted cumulative gain at $K$ between the estimated and oracle strengths satisfies
$\mathrm{NDCG@}K(\hat{S},S^{\star})\;\longrightarrow\;1\quad\text{as }n\to\infty,\;L\to\infty.$
In particular, the NDCG@5 metric used in our experiments is a consistent measure of the ability of ORACLE to recover the top-$K$ target interaction strengths.

### v4

@@ Abstract
We introduce ORACLE, a framework for explaining neural networks on tabular data and scientific factorial designs. ORACLE summarizes a trained network’s prediction surface with main effects and pairwise interactions by treating the network as a black-box response, discretizing the inputs onto a grid, and fitting an orthogonal factorial (ANOVA-style) surrogate—the $L^{2}$ orthogonal projection of the model response onto a finite-dimensional factorial subspace. A simple centering and $\mu$-rebalancing step then expresses this surrogate as main- and interaction-effect tables that remain faithful to the original model in the $L^{2}$ sense. The resulting grid-based interaction maps are easy to visualize, comparable across backbones, and directly aligned with classical design-of-experiments practice. On synthetic factorial benchmarks and low- to medium-dimensional tabular regression tasks, ORACLE more accurately recovers ground-truth interaction structure and hotspots than Monte Carlo SHAP-family interaction methods, as measured by ranking, localization, and cross-backbone stability. We also discuss its scope in latent image and text settings: grid-based factorial surrogates are most effective when features admit an interpretable factorial structure, making ORACLE particularly well-suited to scientific and engineering workflows that require stable DoE-style interaction summaries.
@@ 2 Related Work
Most global interaction explanations build on Shapley interactions. Faith-Shap proposes an axiomatic interaction index (Tsai et al., 2023), and SHAP-IQ/SVARM-IQ/KernelSHAP-IQ provide practical estimators for any-order interactions (Fumagalli et al., 2023; Kolpaczki et al., 2024; Fumagalli et al., 2024), consolidated in shapiq (Muschalik et al., 2024). Related analyses connect global effects to main/interaction decompositions (Herbinger et al., 2024) and apply Shapley interactions to structured phenomena in large models (Singhvi et al., 2024). These approaches typically output scalar indices (not interaction maps) and are rarely evaluated against an oracle or for stability across independently trained backbones.
@@
Functional ANOVA has motivated additive and low-order-interaction models and neural architectures that parameterize ANOVA components or screen interactions with guarantees (Hu et al., 2025; Köhler et al., 2025; Choi et al., 2025; Park et al., 2025). In engineering design and sensitivity analysis, ANOVA-based indices and interpretable networks are used to study physical interactions from observational data (Dolar et al., 2024), and DoE work couples (fractional) factorial designs with ML to estimate effects efficiently (Fontana et al., 2023; Zubair et al., 2024). We adopt an ANOVA/DoE viewpoint but explain a fixed neural backbone via pairwise interaction maps and benchmark interaction explanation methods, rather than proposing another globally additive predictive model.
@@
A complementary line of work develops 2D effect visualizations such as ALE/PDP-style regional plots and fast pipelines (e.g., REPID/DALE) (Herbinger et al., 2022; Gkolemis et al., 2023, 2024a). Systems such as vivid, PDPilot, and Effector support heatmaps and interactive exploration of 1D/2D patterns (Inglis et al., 2022; Kerrigan et al., 2025; Gkolemis et al., 2024b), and transformed-feature-space approaches provide alternatives under strong dependence (Brenning, 2023). Our interaction maps are factorial surrogates with explicit main and pairwise tables, designed for ranking and benchmarking rather than solely visualization.
@@
Prior work studies stability of explanations under perturbations or retraining, including formal metrics and smoothing-based guarantees (Agarwal et al., 2022; Xue et al., 2023), as well as objectives and generative schemes that regularize stability (Chen et al., 2024; Xiang et al., 2023). Other studies report substantial disagreement across independently trained models and propose consistency metrics under distributional or adversarial variation (Watson et al., 2022; Wu et al., 2024; Rai et al., 2024). In contrast, we target the consistency of global pairwise interaction rankings and interaction maps across independently trained backbones, measured via cross-backbone ranking and surface-alignment metrics.
@@
Several works evaluate interaction explanations against controlled structure or discovery guarantees, including formal-language dependencies, Rashomon-set analyses, and error-controlled procedures (Jumelet and Zuidema, 2023; Li et al., 2023; Chen et al., 2025). Recent benchmark efforts propose unit-test-style suites and LLM-scale setups with synthetic triggers (Lee et al., 2025; Kang et al., 2025), and unified evaluation frameworks have been proposed in NLP (Sun et al., 2025). We instead develop a targeted protocol and metrics for global pairwise interaction maps in regression backbones, including oracle recovery and cross-backbone transfer.
@@ 3.2 Target: interaction map of a backbone
$g_{jk}(x_{j},x_{k})$ $=\mathbb{E}\!\left[f(X)\mid X_{j}=x_{j},\,X_{k}=x_{k}\right]$
$\quad-m_{j}(x_{j})-m_{k}(x_{k})-\mu$
@@ 3.3 Discretized factorial surrogate
We discretize each feature into $L$ bins, inducing discretized levels and an implicit $L^{d}$ grid over $\mathcal{X}$. For feature $j$, let $\mathcal{B}_{j}=\{1,\dots,L\}$ index the bins, and let $q_{j}:\mathcal{X}_{j}\to\mathcal{B}_{j}$ map a value $x_{j}$ to its bin index (e.g., via empirical quantiles or uniform spacing over the observed range). A discretized input is represented by a level vector $\ell=(\ell_{1},\dots,\ell_{d})$ with $\ell_{j}\in\mathcal{B}_{j}$.
On these levels, we define a linear factorial surrogate
$\widehat{f}_{L}(\ell)\;=\;\widehat{\mu}+\sum_{j=1}^{d}\widehat{m}_{j}(\ell_{j})+\sum_{1\leq j<k\leq d}\widehat{g}_{jk}(\ell_{j},\ell_{k}),$
where $\widehat{m}_{j}:\mathcal{B}_{j}\to\mathbb{R}$ and $\widehat{g}_{jk}:\mathcal{B}_{j}\times\mathcal{B}_{k}\to\mathbb{R}$ are discrete main-effect and pairwise interaction tables.
We parameterize the surrogate as $\widehat{f}_{L}(\ell)=\langle\widehat{\beta},\varphi(\ell)\rangle$ for a factorial basis vector $\varphi(\ell)\in\mathbb{R}^{p}$ (equivalently, a factorial design matrix with indicator/contrast-coded columns). Importantly, we never enumerate all $L^{d}$ grid points. Instead, we fit by least squares on the sample $\{(x_{i},z_{i})\}_{i=1}^{n}$, where $z_{i}:=q(x_{i})=(q_{1}(x_{i1}),\dots,q_{d}(x_{id}))$. Specifically, ORACLE solves
$\widehat{\beta}\;=\;\arg\min_{\beta}\frac{1}{n}\sum_{i=1}^{n}\Bigl(f(x_{i})-\langle\beta,\varphi(z_{i})\rangle\Bigr)^{2},$
where $\varphi(z_{i})$ denotes the basis evaluated at the discretized levels. The fitted coefficients $\widehat{\beta}$ can be reshaped into $\widehat{\mu}$, $\{\widehat{m}_{j}\}_{j=1}^{d}$, and $\{\widehat{g}_{jk}\}_{1\leq j<k\leq d}$.
@@ 3.4 Recentering and $\mu$-rebalancing (identifiability)
A naive least-squares fit of a factorial surrogate does not necessarily satisfy the standard identifiability constraints for effect tables, such as zero-mean main effects and zero row/column sums for interaction tables. To obtain a well-defined and comparable decomposition, we apply a lightweight post-processing step that recenters the estimated tables and then rebalances the overall mean.
Recentering main effects.
For each feature $j$, we recenter the main-effect table by subtracting its average so that
$\sum_{\ell_{j}\in\mathcal{B}_{j}}\widehat{m}_{j}(\ell_{j})\;=\;0.$
Recentering interactions.
For each pair $(j,k)$, we recenter the interaction table $\widehat{g}_{jk}$ so that both its row and column sums vanish:
$\sum_{\ell_{j}\in\mathcal{B}_{j}}\widehat{g}_{jk}(\ell_{j},\ell_{k})$ $=0,\qquad$ $\text{for all }\ell_{k}\in\mathcal{B}_{k},$
$\sum_{\ell_{k}\in\mathcal{B}_{k}}\widehat{g}_{jk}(\ell_{j},\ell_{k})$ $=0,\qquad$ $\text{for all }\ell_{j}\in\mathcal{B}_{j}.$
$\mu$-rebalancing.
The centering operations may induce an additive constant shift. We absorb this offset into $\widehat{\mu}$ so that the fitted surrogate function itself is preserved. After recentering and $\mu$-rebalancing, the surrogate admits the decomposition
$\widehat{f}_{L}(\ell)\;=\;\widehat{\mu}+\sum_{j=1}^{d}\widehat{m}_{j}(\ell_{j})+\sum_{1\leq j<k\leq d}\widehat{g}_{jk}(\ell_{j},\ell_{k}),$
with the usual sum-to-zero constraints at the discretized level.
These constraints are identifiability conventions: they do not change the fitted surrogate $\widehat{f}_{L}$, but they fix a unique representation in terms of main and pairwise effect tables. Moreover, they mirror the conditional-centering constraints that characterize orthogonal (discrete) ANOVA components under the discretized distribution $P_{Z}$ (see Appendix A.2A.3). We summarize the recentering and $\mu$-rebalancing procedure in Algorithm 1 (Appendix B).
Finally, we interpret the resulting tables as discrete approximations to the corresponding functional-ANOVA components of $f$ under $P_{X}$; the formal connection via an $L^{2}(P_{Z})$ projection is given in Appendix A.3.
@@ 3.5 Extracting interaction maps and strengths
The post-processed interaction tables $\widehat{g}_{jk}:\mathcal{B}_{j}\times\mathcal{B}_{k}\to\mathbb{R}$ define discrete interaction maps for each feature pair $(j,k)$. These maps can be visualized as heatmaps over the discretized levels $(\ell_{j},\ell_{k})\in\mathcal{B}_{j}\times\mathcal{B}_{k}$ (and, by extension, over the corresponding regions of $(X_{j},X_{k})$), revealing where the backbone exhibits strong non-additivity.
To summarize interaction strength in a single scalar, we compute an RMS with respect to the empirical distribution of the discretized pair $(Z_{j},Z_{k})$:
$\widehat{S}_{jk}$ $=\Bigg(\sum_{\ell_{j}\in\mathcal{B}_{j}}\sum_{\ell_{k}\in\mathcal{B}_{k}}\widehat{p}_{jk}(\ell_{j},\ell_{k})\,\widehat{g}_{jk}(\ell_{j},\ell_{k})^{2}\Bigg)^{1/2},$
$\widehat{p}_{jk}(\ell_{j},\ell_{k})$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{z_{ij}=\ell_{j},\,z_{ik}=\ell_{k}\}.$
where $z_{i}:=q(x_{i})$ and $z_{ij}$ denotes its $j$th component. We use $\{\widehat{S}_{jk}\}_{j<k}$ to rank interactions.
In our experiments, the collection of interaction maps $\{\widehat{g}_{jk}\}_{j<k}$ and strengths $\{\widehat{S}_{jk}\}_{j<k}$ constitutes ORACLE’s global explanation. We evaluate them against oracle interaction maps derived from the backbone under $P_{X}$, and compare to SHAP-family baselines in terms of interaction selection, hotspot localization, cross-backbone transfer, scale alignment, and intervention-oriented utility.
@@ 4.1 Oracle interaction decomposition via functional ANOVA
$g_{jk}(x_{j},x_{k})$ $=\mathbb{E}\!\left[f(X)\mid X_{j}=x_{j},\,X_{k}=x_{k}\right]$
$\quad-m_{j}(x_{j})-m_{k}(x_{k})-\mu.$
and this representation coincides with the unique orthogonal projection of $f$ onto $H_{\leq 2}$ (see Proposition A.5 in Appendix A).
@@ 4.3 Consistency of interaction strengths and Top-$K$ selection
We study the asymptotic behavior of interaction strengths induced by the discretized interaction maps. Fix a grid size $L$ and recall the discretized features $Z_{j}=q_{j}(X_{j})\in\mathcal{B}_{j}$. Define the discrete interaction strength
$S^{L}_{jk}=\Big(\mathbb{E}\big[\big(g^{L}_{jk}(Z_{j},Z_{k})\big)^{2}\big]\Big)^{1/2},$
where the expectation is with respect to the distribution of $(Z_{j},Z_{k})$ induced by $P_{X}$. We estimate $S^{L}_{jk}$ by the empirical (plug-in) counterpart
$\widehat{S}_{jk}$ $=\Big(\mathbb{E}_{n}\big[\big(\widehat{g}_{jk}(Z_{j},Z_{k})\big)^{2}\big]\Big)^{1/2}$
$=\Bigg(\sum_{\ell_{j}\in\mathcal{B}_{j}}\sum_{\ell_{k}\in\mathcal{B}_{k}}\widehat{p}_{jk}(\ell_{j},\ell_{k})\,\widehat{g}_{jk}(\ell_{j},\ell_{k})^{2}\Bigg)^{1/2},$
where $\widehat{p}_{jk}(\ell_{j},\ell_{k})=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{z_{ij}=\ell_{j},\,z_{ik}=\ell_{k}\}$ and $z_{i}:=q(x_{i})$.
@@
Under the conditions of Proposition 4.6, for any fixed $L$,
$\widehat{S}_{jk}\xrightarrow{p}S^{L}_{jk}\quad\text{for all }1\leq j<k\leq d\quad\text{as }n\to\infty.$
If, in addition, the oracle interaction maps $g_{jk}$ are square-integrable and sufficiently regular, then for each $(j,k)$,
$S^{L}_{jk}\to S^{\star}_{jk}\quad\text{as }L\to\infty,$
where $S^{\star}_{jk}$ denotes the corresponding oracle interaction strength.
We next translate these convergences into guarantees for ranking and NDCG-style metrics. Let $S^{\star}_{(1)}\geq\dots\geq S^{\star}_{(M)}$ denote the sorted oracle strengths over all $M=\binom{d}{2}$ pairs, and let $\widehat{S}_{(1)}\geq\dots\geq\widehat{S}_{(M)}$ be the sorted estimated strengths.
@@
for some fixed $K\in\{1,\dots,M-1\}$. Then,
$\lim_{L\to\infty}\;\lim_{n\to\infty}\;\mathbb{P}\!\left(\mathrm{Top}\text{-}K(\widehat{S})=\mathrm{Top}\text{-}K(S^{\star})\right)=1,$
where $\mathrm{Top}\text{-}K(\cdot)$ returns the set of indices of the $K$ largest entries.
@@
Assume the conditions of Theorem 4.8 and, in addition, that the oracle strengths within the top-$K$ are separated: there exists $\Delta_{\mathrm{in}}>0$ such that
$S^{\star}_{(r)}-S^{\star}_{(r+1)}\geq\Delta_{\mathrm{in}}\qquad\text{for all }r\in\{1,\dots,K-1\}.$
Then,
$\lim_{L\to\infty}\;\lim_{n\to\infty}\;\mathrm{NDCG@}K(\widehat{S},S^{\star})=1.$
In particular, the NDCG@5 metric used in our experiments is a consistent measure of the ability of ORACLE to recover the top-$K$ interaction strengths and their ordering.